# Scalable Graph Generative Modeling via Substructure Sequences

### v1

@@ 7 Conclusion
We introduce G2PM, a Transformer-based generative pretraining framework for graphs that operates entirely without message passing. Unlike traditional GNNs, G2PM scales effectively with both data and model size, and consistently improves performance across tasks. Through extensive ablations and experiments, we demonstrate its expressiveness, transferability, and potential as a scalable backbone for graph representation learning. Limitation and Future Works: While G2PM leverages unordered substructure sequences—suitable for masked-token prediction—extending it to ordered sequences may enable next-token prediction, further improving scalability. Additionally, our use of random walks as an online tokenizer opens up future directions in designing learnable and adaptive substructure tokenizers for graphs.
@@ Appendix A From the Perspective of Substructure Dependencies
Functional Reinforcement. Certain substructures statistically co-occur due to the generative processes underlying different graph types [43]. For instance, in social and citation networks, a high density of triangles often correlates with the existence of higher-order cliques [6, 4]. Similarly, the frequent appearance of short chains may indicate the potential for longer cycle [31, 45]. We refer to this pattern as functional reinforcement, where the presence of one motif increases the likelihood of encountering another. This mutual reinforcement reflects domain structural priors and can be leveraged during pre-training to build a predictive inductive bias across motifs.
Functional Exclusion. Conversely, some substructures exhibit negative correlations, reflecting competition for structural space within the graph. We term this phenomenon functional exclusion. For example, a node embedded within multiple triangle-like motifs is likely part of a densely connected community, making it less probable to support hub-like star patterns [3, 53]. Likewise, graphs with tree-like branches typically lack the density required to support large cliques [1]. Understanding such exclusion enables the model to refine its prediction space: it can down-weight structurally incompatible or redundant patterns during reconstruction.

### v3

@@ 7 Conclusion
We introduce G2PM, a Transformer-based generative pretraining framework for graphs that operates entirely without message passing. Unlike traditional GNNs, G2PM scales effectively with both data and model size, and consistently improves performance across tasks. Through extensive ablations and experiments, we demonstrate its expressiveness, transferability, and potential as a scalable backbone for graph representation learning. While G2PM leverages unordered substructure sequences—suitable for masked-token prediction—extending it to ordered sequences may enable next-token prediction, further improving scalability. Additionally, our use of random walks as an online tokenizer opens up future directions in designing learnable and adaptive substructure tokenizers for graphs.
Acknowledgement
The work was partially supported by the NSF under grants IIS-2533550, IIS-2321504, IIS-2340346, IIS-2217239, IIS-2528540, CNS-2426514, and CMMI-2146076, ND-IBM Tech Ethics Lab Program, Notre Dame Strategic Framework Research Grant (2025), and Notre Dame Poverty Research Package (2025). Any expressed opinions, findings, and conclusions or recommendations are those of the authors and do not necessarily reflect the views of the sponsors.
@@ Appendix A From the Perspective of Substructure Dependencies
Functional Reinforcement. Certain substructures statistically co-occur due to the generative processes underlying different graph types [ruiz2020graphon]. For instance, in social and citation networks, a high density of triangles often correlates with the existence of higher-order cliques [bianconi2014triadic, asikainen2020cumulative]. Similarly, the frequent appearance of short chains may indicate the potential for longer cycle [li2021analyzing, schwarze2021motifs]. We refer to this pattern as functional reinforcement, where the presence of one motif increases the likelihood of encountering another. This mutual reinforcement reflects domain structural priors and can be leveraged during pre-training to build a predictive inductive bias across motifs.
Functional Exclusion. Conversely, some substructures exhibit negative correlations, reflecting competition for structural space within the graph. We term this phenomenon functional exclusion. For example, a node embedded within multiple triangle-like motifs is likely part of a densely connected community, making it less probable to support hub-like star patterns [arenas2008motif, tsourakakis2017scalable]. Likewise, graphs with tree-like branches typically lack the density required to support large cliques [abu2016metric]. Understanding such exclusion enables the model to refine its prediction space: it can down-weight structurally incompatible or redundant patterns during reconstruction.